OpenStudy (anonymous):

write the equation in the point slope form (-2,3);m=-2

OpenStudy (anonymous):

Do you know what point slope form is?

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

Correct. So since you know the slope and a point, you can plug in.

OpenStudy (anonymous):

the point (-2, 3) represents (x1, y1)

OpenStudy (anonymous):

y-3=m(x-2)

OpenStudy (anonymous):

No but Close, be careful with your signs. Check it again.

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

y=3/2 -2

OpenStudy (anonymous):

Umm, no.

OpenStudy (anonymous):

Look below: